A. 200 raffel tickets are sold for $1 each. The prize is a $50 gift certificate.   What is the expected value of this game?  Answer as dollars and cents.
Answer #1

For expected value, we have to find the probability of winning and losing

Total Tickets = 200

Win ticket = 1 . and you will win (50-1) = 49 dollars

Losing ticket = 199 and there will be a loss of 1 dollar

P of winning = 1/200

P of losing = 199/200

Expected Value = win * P of win + Loss * P of loss

= 49* (1/200) - 1 * (199/200) = -0.75

ans. -0.75 dollars = - 75 cents

Loss of 75 cents